A baseball is hit vertically at a velocity of 50 ft per sec. What is the maximum height that the ball will reach, and after how many seconds does the ball reach its maximum height? (The height after t seconds is given by h=50t - 16t^2)
Answer by stanbon(57940)

A baseball is hit vertically at a velocity of 50 ft per sec. What is the maximum height that the ball will reach, and after how many seconds does the ball reach its maximum height? (The height after t seconds is given by
h=50t - 16t^2)
--------
Maximum height occurs when t = -b/(2a) = -50/(2*-16) = 1.56 seconds
-----
Max height = h(1.56) = 39.06 ft.
=================
Cheers,
Stan H.
=================